# Channel-Selective Normalization for Label-Shift Robust Test-Time Adaptation

### v1

@@ 4 Experiments and Results
We use three datasets in our evaluations: two popular benchmarks, CIFAR-10-C and ImageNet-1K-C, and a publicly available fatty liver ultrasound dataset [6].
Figure 2: On CIFAR-10 we adapt models only up to the layer shown on the x-axis, the y-axis showing the performance on the target data. We consider label distributions with all (10) classes as well as 5,3, and 1 randomly selected and balanced classes. We observe that adapting later layers has an outsized role in the catastrophic collapse due to TTN
@@
CIFAR-10 and CIFAR-10-C. We use the CIFAR-10 [20] dataset along with CIFAR-10-C [15]. CIFAR-10 is a small natural image dataset with 50k training images and 10k validation images. CIFAR-10-C contains corrupted versions of the CIFAR-10 Validation set at varying severities. We train our models on the uncorrupted dataset.
@@
ImageNet-1K and ImageNet-1K-C. We use the ImageNet-1K [10] dataset along with ImageNet-1K-C [15]. ImageNet-1K is a large natural image dataset with 1.2 million training images and 50k validation images. ImageNet-1K-C, similarly to CIFAR-10-C, contains corrupted versions of the ImageNet-1K Validation set at varying severities. Both CIFAR-10-C and ImageNet-1K-C are popular as a measure of robustness to covariate shift.
@@
Liver Ultrasound. We use, as the training data, a dataset of ultrasound abdominal images collected at a liver transplantation center from patients under suspicion of fatty liver disease [42]. Images were collected between September 2011 to October 2020 with seven different devices, with varying settings, and by multiple operators. It contains 7,529 B-mode images from 57 control cases without NAFLD and 142 patients with NAFLD. We train our models for binary classification, i.e. detecting fatty liver versus non-fatty liver, on this dataset. For the test data in TTN, we use a publicly available dataset collected at a different institution [6], with 550 images. The differences between the source and the target data are multiple. On the patients level, we have differences in population demographics and in the inclusion criteria, e.g., the 55 patients on the publicly available dataset are diagnosed with obesity, which impacts the wave propagation during ultrasound exams. On the examinations level, the second dataset was collected prospectively, with consistent settings and using a single ultrasound machine, which was not present in the source data collection. This creates a more homogeneous dataset and provides the covariate shift for the TTN method. Finally, the training dataset has 84% of its images from cases of fatty liver, while the test dataset has 69%, introducing the label distribution shift. For both datasets, all patients underwent liver biopsy and the steatosis grades in the pathology reports are used as the reference standard for diagnosis of fatty liver. In order to introduce higher levels of covariate shift akin to CIFAR-10 and ImageNet-1K evaluations, we added speckle noise to the target dataset, with varying levels of corruption.
@@
Architecture Details
On CIFAR-10 we train a Resnet-26 model as defined in [13]. For ImageNet-1K we use a pre-trained Resnet18 model. For the Liver Ultrasound data, we train a Resnet-18 model.
Adaptation Details We focus on the source-free domain adaptation setting where adaptation is done on a single batch without affecting the deployed model [26]. For CIFAR-10-C and ImageNet-1K-C we use a batch size of 500 for the experiments (sampled over multiple seeds). For Liver Ultrasound, the batch size is the length of the target set. For the Hybrid-TTN, the percentage of not-adapted channels by layer is defined by a linear function starting at 0% at the first BatchNorm layer and ending at 100% at the last layer. The step of this linear function changes by model (e.g., Resnet-26 has more BatchNorm layers than Resnet-18), but it does not introduce any hyperparameters. Our method is hyperparameter free and appplied on all datasets and settings in the same way.
Target Label Distributions In order to analyze label distribution shifts we consider two approaches for constructing class imbalanced target distributions, while we keep the source training balanced. $N$-class refers to randomly selecting $N$ classes from the set of all classes and constructing a batch with the specified $N$ classes. We also use a more natural imbalanced dataset simulation construction technique popular in the federated learning literature [23] which samples from a dirichlet distribution with parameter $\alpha$. Smaller $\alpha$ will yield more sample batches with higher imbalances. This allows us to mimic a more natural scenario where the label distribution is not only varying in classes but proportions. In Liver Ultrasound, we use the original target distribution, which is different from the original source distribution, alongside a balanced distribution, and severely imbalanced cases.
@@
4.1 Label Distribution Shift Impact on TTN
Figure 5: CIFAR-10 evaluations on multiple label shifted distributions and covariate shifts (corruptions) with different degrees of label imbalance. We show the source model accuracy and the improvement (or degradation) with our proposed method and baseline. We observe that the proposed method provides benefits when there is no covariate shift, while avoiding catastrophic failures and allowing benefits over source when there are label distribution shifts.
We first illustrate the potential pitfalls of TTN on several examples of label distribution shift. Using the CIFAR-10 dataset we show the effect on TTN of an extreme version of label distribution shift in Figure 1. Specifically, we adapt on batches containing one class or all classes (as can be seen in the label distributions shown in Figure 1). We observe in the adaptation of one class batches that the accuracy of TTN degrades severely compared to the source model in cases where there is a covariate shift (application of Gaussian noise to the input) but also in the absence of covariate shift. The latter observation is also confirmed in the Appendix for ultrasound data where we modify the target label proportion of the test data coming from the same distribution (but without adding noise). This catastrophic failure, even without covariate shift, is clearly undesired when operating in a realistic deployment environment where the test set label distribution is unknown and can vary for each incoming batch.
Furthermore, we can observe from the per-class performance evaluations of the adapted models that the majority class is disproportionately affected by TTN, and that classes underrepresented (or not present) will, unintuitively, experience less degradation or even slightly improved performance compared to source model. We can obtain intuition about this latter behavior by considering a strong classification model where representations from each class will be well separated and clustered in the final representation layer (see Figure 3). Note that this is an assumption that is correct for most modern neural network architectures, as it is a consequence of the neural collapse phenomenon first described in [32]. A biased BatchNorm computation will compute the mean of a specific label cluster and re-center the data by this mean. Now, consider a neural network layer whose activations are tightly clustered, such that each cluster corresponds to one label. Further, assume that the different cluster centers occupy approximately orthogonal directions in this space. As mentioned above, such quasi-orthogonality is not an uncommon property in high dimensions [32]. In this setting, linear decision directions would align with cluster means. Subtracting a cluster mean instead of a class-balanced batch mean would then significantly move the label cluster along its decision direction and will likely lead to some examples passing below the classification threshold. Other classes will be barely affected due to the orthogonality of the cluster centers and hence of the decision boundaries.
Figure 6: ImageNet-1K evaluations on multiple label shifted distributions and covariate shifts (corruptions) with different degrees of label imbalance. We observe that the proposed method provides benefits when there is covariate shift, while avoiding catastrophic failures when there are label distribution shifts.
4.2 Label Sensitive Channel Level Adaptation
As mentioned previously, many works in deep learning have shown that supervised trained neural networks become more specialized as depth of a network increases [52, 2, 31] and that classes progressively become more separated with depth. This suggests the later layers are more likely to have channels which are very sensitive to specific classes.
Adaptation of Early Layers
To further confirm that depth can have an effect on the behavior of TTN and taking into account that earlier layers tend to be less specialized, we now consider only adapting earlier parts of the network and study this effect on the classification performance under label distribution shift. We perform experiments on layer-limited adaptation for a Resnet-26 model on CIFAR-10 both with and without noise. Our results are shown in in Figure 2. Specifically, for each index on the x-axis we perform test-time normalization of the model up to this layer. Consequently, the extremes of the x-axis (0 and, here, 30) correspond to Source and TTN models respectively. First, we observe that on the non-corrupted train data the performance of class-imbalanced data degrades gradually at first and increasingly faster towards the later layers. This suggests that later layers can cause a large degradation. Secondly, for corrupted data we observe that adapting up to earlier layers can allow enough label distribution invariance to provide benefits under covariate shift. For example, models that are adapted up to layers 8 through 15 can perform much better than source models under covariate shift but without label distribution shift, while also being able to perform much better than fully adapted and often source models in various label distribution shift scenarios. This suggests limiting adaptation in later layers can be beneficial. On the other hand, it is not obvious a priori which layer should be adapted. Furthermore, different classes can have more fine-grained effects, for example, some classes may lead to over-sensitivity in certain neurons or channels in a layer than others.
@@ 4.2 Label Sensitive Channel Level Adaptation
Figure 7: Evaluation on ultrasound data. Binary classification, subjects with and without fatty liver disease. Models were trained on one site and are evaluated on a publicly available dataset ([6]). We show evaluations on targets with shifts in label distribution. We observe that the proposed method is able to control the degradation in performance under severe label distribution shift.
@@ 4.3 Evaluating Hybrid-TTN
We now use Hybrid-TTN method on a variety of target datasets, covariate shifts, and label distribution shifts. We demonstrate that Hybrid-TTN can provide a good trade-off in being able to adapt to covariate shift without experiencing catastrophic failure due to label distribution shift. Our results are shown in Figures 5, 6, and 7, and detailed tables are available in the Appendix. Here we demonstrate various degrees of covariate and label distribution shift, and the gain or loss as compared to the source model performance.
@@
Table 2: Ablations. We use CIFAR-10 with batches from a single class and different degrees of covariate shift. We show that our Wassertstein scoring and class prior estimates are both beneficial to Hybrid-TTN.
4.4 Ablation of Channel Selection
We perform an ablation study aimed at investigating the efficacy of the different elements of the channel selection strategy within Hybrid-TTN.
Class Prior We now consider the importance of the pseudo-labeling step used to obtain a class prior from the target data described in Section 3. Specifically, we evaluate the performance of just using a uniform prior, we observe this improves the performance from the random channel scores, but substantially underperforms Hybrid-TTN. Finally we also show the performance of an oracle class prior, where the exact target data label distribution is given. We note that for none or low corruption cases this gives performance comparable to Hybrid-TTN, further affirming our approach. As the corruption becomes more severe the performance gap between the oracle and our method increases.
@@ 5 Conclusions
We have studied a popular batch-level Test-time Adaptation method in the context of label distribution shift. We observed that in realistic scenarios where batches at deployment time have label distribution shifts, this method can fail catastrophically. We proposed a direction for solving this problem to keep the benefits of adaptation without risking catastrophic failure due to label shift. Moreover, our results indicate that the challenge of domain adaptation in the biomedical field, specifically in diagnosis of fatty liver using grayscale ultrasound images, can be addressed by the proposed Hybrid-TTN method, reducing risks in the deployment of a trained framework in different populations.
Supplementary Material
A Appendix
@@ A.1 Training details
On CIFAR-10 we train a Resnet-26 model as defined in [13]. We use an SGD optimizer with a batch size of 128. An initial learning rate set to 0.1 is used in combination with a cosine annealing schedule [29] trained over 200 epochs. Weight decay set to 5e-4 is used along with momentum set to 0.9 [51]. Standard augmentation uses random crop of size 32 with 4 padding, and random horizontal flips.
@@ A.3 Example of failure case in Ultrasound
A.4 Detailed results
Below, we present the tables for all of our experiments.

### v2

@@ 4 Experiments and Results
In this section we discuss our experimental setup. We use three datasets in our evaluations: two popular benchmarks, CIFAR-10-C and ImageNet-1K-C, and a publicly available fatty liver ultrasound dataset.
@@
CIFAR-10 and CIFAR-10-C. We use the CIFAR-10 (Krizhevsky, 2009) dataset along with CIFAR-10-C (Hendrycks & Dietterich, 2019b). CIFAR-10 is a small natural image dataset with 50k training images and 10k validation images. CIFAR-10-C contains corrupted versions of the CIFAR-10 Validation set at varying severities. We train our models on the uncorrupted dataset.
@@
ImageNet-1K and ImageNet-1K-C. We use the ImageNet-1K (Deng et al., 2009a) dataset along with ImageNet-1K-C (Hendrycks & Dietterich, 2019b). ImageNet-1K is a large natural image dataset with 1.2 million training images and 50k validation images. ImageNet-1K-C, similarly to CIFAR-10-C, contains corrupted versions of the ImageNet-1K Validation set at varying severities. Both CIFAR-10-C and ImageNet-1K-C are popular as a measure of robustness to covariate shift.
@@
Liver Ultrasound. We use, as the training data, a dataset of ultrasound abdominal images collected at a liver transplantation center from patients under suspicion of non-alcoholic fatty liver disease (NAFLD) (Vianna et al., 2023). Images were collected between September 2011 to October 2020 with seven different devices, with varying settings, and by multiple operators. It contains 7,529 B-mode images from 57 control cases without NAFLD and 142 patients with NAFLD. We train our models for binary classification, i.e. detecting fatty liver versus non-fatty liver, on this dataset. For the test data in TTN, we use a publicly available dataset collected at a different institution (Byra et al., 2018), with 550 images. The differences between the source and the target data are multiple. On the patients level, we have differences in population demographics and in the inclusion criteria, e.g., the 55 patients on the publicly available dataset are diagnosed with obesity, which impacts the wave propagation during ultrasound exams. On the examinations level, the second dataset was collected prospectively, with consistent settings and using a single ultrasound machine, which was not present in the source data collection. This creates a more homogeneous dataset and provides the covariate shift for the TTN method. Finally, the training dataset has 84% of its images from cases of fatty liver, while the test dataset has 69%, introducing the label distribution shift. For both datasets, all patients underwent liver biopsy and the steatosis grades in the pathology reports are used as the reference standard for diagnosis of fatty liver. In order to introduce higher levels of covariate shift akin to CIFAR-10 and ImageNet-1K evaluations, we added speckle noise to the target dataset, with varying levels of corruption.
Architecture Details On CIFAR-10 we train a Resnet-26 model as defined in (He et al., 2016). For ImageNet-1K we use a pre-trained Resnet-18 model. For the Liver Ultrasound data, we train a Resnet-18 model. Training details are described on Appendix A.1.
Adaptation Details We focus on the source-free domain adaptation setting where adaptation is done offline on a single batch without affecting the deployed model (Liang et al., 2023), avoiding issues with catastrophic forgetting (Niu et al., 2022). For CIFAR-10-C and ImageNet-1K-C we use a batch size of 500 for the experiments (sampled over multiple seeds). For Liver Ultrasound, the batch size is the length of the target set. For the Hybrid-TTN, the percentage of not-adapted channels by layer is defined by a linear function starting at 0% at the first BatchNorm layer and ending at 100% at the last layer. The step of this linear function changes by model (e.g., Resnet-26 has more BatchNorm layers than Resnet-18), but it does not introduce any hyperparameters. Our method is hyperparameter free and appplied on all datasets and settings in the same way. More adaptation details are described in the Appendix A.2.
Target Label Distributions In order to analyze label distribution shifts we consider two approaches for constructing class imbalanced target distributions, while we keep the source training balanced. The first approach is to randomly select $N$ classes from the set of all classes and constructing a batch with the specified $N$ classes. We also use a more natural imbalanced dataset simulation construction technique popular in the federated learning literature which samples from a Dirichlet distribution with parameter $\alpha$ (Li et al., 2021). Smaller $\alpha$ will yield more sample batches with higher imbalances. This allows us to mimic a more natural scenario where the label distribution is not only varying in classes but proportions. In Liver Ultrasound, we use the original target distribution, which is different from the original source distribution, alongside a balanced distribution, and severely imbalanced cases.
4.1 Label Distribution Shift Impact on TTN
We first illustrate the potential pitfalls of TTN on several examples of label distribution shift. Using the CIFAR-10 dataset we show the effect on TTN of an extreme version of label distribution shift in Figure 1. Specifically, we adapt on batches containing one class or all classes (as can be seen in the label distributions shown in Figure 1). We observe in the adaptation of one class batches that the accuracy of TTN degrades severely compared to the source model in cases where there is a covariate shift (application of Gaussian noise to the input) but also in the absence of covariate shift. The latter observation is also confirmed for ultrasound data where we modify the target label proportion of the test data coming from the same distribution, but without adding noise (in Appendix A.3). This catastrophic failure, even without covariate shift, is clearly undesired when operating in a realistic deployment environment where the test set label distribution is unknown and can vary for each incoming batch.
Furthermore, we can observe from the per-class performance evaluations of the adapted models that the majority class is disproportionately affected by TTN, and that classes underrepresented (or not present) will, unintuitively, experience less degradation or even slightly improved performance compared to source model. We can obtain intuition about this latter behavior by considering a strong classification model where representations from each class will be well separated and clustered in the final representation layer (see Figure 2). Note that this is an assumption that is correct for most modern neural network architectures, as it is a consequence of the neural collapse phenomenon first described in Papyan et al. (2020). A biased BatchNorm computation will compute the mean of a specific label cluster and re-center the data by this mean. Now, consider a neural network layer whose activations are tightly clustered, such that each cluster corresponds to one label. Further, assume that the different cluster centers occupy approximately orthogonal directions in this space. As mentioned above, such quasi-orthogonality is not an uncommon property in high dimensions (Papyan et al., 2020). In this setting, linear decision directions would align with cluster means. Subtracting a cluster mean instead of a class-balanced batch mean would then significantly move the label cluster along its decision direction and will likely lead to some examples passing below the classification threshold. Other classes will be barely affected due to the orthogonality of the cluster centers and hence of the decision boundaries.
@@ 4.1 Label Distribution Shift Impact on TTN
4.2 Label Sensitive Channel Level Adaptation
As mentioned previously, many works in deep learning have shown that supervised trained neural networks become more specialized as depth of a network increases (Zeiler & Fergus, 2014; Belilovsky et al., 2019; Oyallon, 2017) and that classes progressively become more separated with depth. This suggests the later layers are more likely to have channels which are very sensitive to specific classes.
Adaptation of Early Layers To further confirm that depth can have an effect on the behavior of TTN and taking into account that earlier layers tend to be less specialized, we now consider only adapting earlier parts of the network and study this effect on the classification performance under label distribution shift. We perform experiments on layer-limited adaptation for a Resnet-26 model on CIFAR-10 both with and without noise. Our results are shown in in Figure 3. Specifically, for each index on the x-axis we perform test-time normalization of the model up to this layer. Consequently, the extremes of the x-axis (0 and, here, 30) correspond to Source and TTN models respectively. First, we observe that on the non-corrupted train data the performance of class-imbalanced data degrades gradually at first and increasingly faster towards the later layers. This suggests that later layers can cause a large degradation. Secondly, for corrupted data we observe that adapting up to earlier layers can allow enough label distribution invariance to provide benefits under covariate shift. For example, models that are adapted up to layers 8 through 15 can perform much better than source models under covariate shift but without label distribution shift, while also being able to perform much better than fully adapted and often source models in various label distribution shift scenarios. This suggests limiting adaptation in later layers can be beneficial. On the other hand, it is not obvious a priori which layer should be adapted. Furthermore, different classes can have more fine-grained effects, for example, some classes may lead to over-sensitivity in certain neurons or channels in a layer than others.
Figure 3: On CIFAR-10 we adapt models only up to the layer shown on the x-axis, the y-axis showing the accuracy on the target data. We consider label distributions with all (10) classes as well as 5,3, and 1 randomly selected and balanced classes. We observe that adapting later layers has an outsized role in the catastrophic collapse due to TTN.
@@ 4.3 Evaluating Hybrid-TTN
We now use Hybrid-TTN method on a variety of target datasets, covariate shifts, and label distribution shifts. We demonstrate that Hybrid-TTN can provide a good trade-off in being able to adapt to covariate shift without experiencing catastrophic failure due to label distribution shift. Our results are shown in Figures 5, 6, and 7, and more detailed tables are available in the Appendix A.6 and A.7. Here we demonstrate various degrees of covariate and label distribution shift, and the gain or loss as compared to the source model performance.
Figure 5: CIFAR-10 evaluations on multiple label shifted distributions and covariate shifts (all corruptions) with different degrees of label imbalance. We show the source model accuracy and the improvement (or degradation) with our proposed method and baseline. We observe that the proposed method provides benefits when there is no covariate shift, while avoiding catastrophic failures and allowing benefits over source when there are label distribution shifts.
Figure 6: ImageNet-1K evaluations on multiple label shifted distributions and covariate shifts (all corruptions) with different degrees of label imbalance. We observe that the proposed method provides benefits when there is covariate shift, while avoiding catastrophic failures when there are label distribution shifts.
Figure 7: Evaluation on ultrasound data. Binary classification, subjects with and without fatty liver disease. Models were trained on one site and are evaluated on a publicly available dataset (Byra et al., 2018). We show evaluations on targets with shifts in label distribution. We observe that the proposed method is able to control the degradation in performance under severe label distribution shift.
@@
Computational Complexity In terms of computational complexity, we note that the scores are computed in Algorithm 1 prior to deployment, thus Algorithm 1 is run only once while Algorithm 2 can be used for any new data. Algorithm 2 adds minimal overhead to a standard forward pass. We benchmark our implementation of the Hybrid-TTN approach and note it takes $0.89\pm 0.01$ seconds to output its predictions, compared to $0.39\pm 0.003$ seconds for TTN which is the lowest overhead TTA method. This could be explained by the two passes of Algorithm 2, using a different class prior each time. A comparison of runtimes across different methods evaluated is presented in the Appendix A.4.
4.4 Ablation of Channel Selection
We perform an ablation study aimed at investigating the efficacy of the different elements of the channel selection strategy within Hybrid-TTN. The ablation experiments for a single class across all corruptions is presented in Table 2.
Class Prior We now consider the importance of the pseudo-labeling step used to obtain a class prior from the target data described in Section 3. Specifically, we evaluate the performance of just using a uniform prior, we observe this improves the performance from the random channel scores, but substantially underperforms Hybrid-TTN. Finally we also show the performance of an oracle class prior, where the exact target data label distribution is given. We note that for none or low corruption cases this gives performance comparable to Hybrid-TTN, further affirming our approach. As the corruption becomes more severe the performance gap between the oracle and our method increases. This effect was also observed in additional experiments considering different CNN architectures and CIFAR-100, detailed in the Appendix A.5.
@@ 4.4 Ablation of Channel Selection
Table 2: Results for the ablation study. Each technique delta accuracy is in comparison to the source results. We use CIFAR-10 with batches from a single class and different degrees of covariate shift. We show that our Wassertstein scoring and class prior estimates are both beneficial to Hybrid-TTN.
5 Limitations
A limitation of our proposed technique is that it is designed for architectures featuring batch normalization layers. However, it is noteworthy that this is still a popular kind of architecture in the TTA literature (Zhao et al., 2023). Furthermore, future work can adapt our channel selection criteria to other methods, by using a different kind of scoring function and avoiding adaptation (e.g. to entropy minimization) of selected channels. Another limitation is that we assume access to the post-training but pre-deployment access to the source data, however note that this is practical in most cases except where the entity training the model does not know of the eventual application of TTA to the model.
@@ 6 Conclusions
We have studied a popular batch-level Test-time Adaptation method in the context of label distribution shift. We observed that in realistic scenarios where batches at deployment time have label distribution shifts, this method can fail catastrophically. A comparison in the offline setting indicates that common test-time adaptation techniques are detrimental to classification, contingent upon shifts in label distribution. We proposed a direction for solving this problem to keep the benefits of adaptation without risking catastrophic failure due to label shift. Compared to similar works, our proposed technique was designed to be generic and easily applied without hyperparameter tuning, as we assume a fixed linear depth prior and a uniform distribution prior which are applied to all datasets and architectures studied. Moreover, our results indicate that the challenge of domain adaptation in the biomedical field, specifically in diagnosis of fatty liver using grayscale ultrasound images, can be addressed by the proposed Hybrid-TTN method, reducing risks in the deployment of a trained framework in different populations.
Acknowledgments
This work was supported by grants from the Institute of Data Valorization (IVADO PRF3) to A.T., G.W. and E.B. A.T. acknowledges support from the Fonds de Recherche du Québec – Santé (FRQS), Fonds de Recherche du Québec - Nature et technologies (FRQNT) and the Fondation de l’Association des Radiologistes du Québec (FARQ) Clinical Research Scholarship–Senior Salary Award (FRQS-ARQ no. 298509). G.W. acknowledges support from the Canada CIFAR AI Chair. E.B. and G.W. acknowledge funding from Fonds de recherche du Québec — Nature et technologies - NOVA (2023-NOVA-329759 and 2023-NOVA-329125).
Appendix A Appendix
@@
  • Runtime comparison across techniques (Sec. A.4)
  • Comparison to different architectures and CIFAR-100 (Sec. A.5)
@@ A.1 Training details
On CIFAR-10 we train a Resnet-26 model as defined in He et al. (2016). We use an SGD optimizer with a batch size of 128. An initial learning rate set to 0.1 is used in combination with a cosine annealing schedule (Loshchilov & Hutter, 2016) trained over 200 epochs. Weight decay set to 5e-4 is used along with momentum set to 0.9 (Zagoruyko & Komodakis, 2016). Standard augmentation uses random crop of size 32 with 4 padding, and random horizontal flips.
@@ A.3 Example of failure case in Ultrasound
A.4 Adaptation runtimes
Below we present the table for the runtime of each investigated test-time adaptation technique. Values are averaged over 50 random seeds and each experiment is for the process time of a batch of size 500 of CIFAR-10 images.
Algorithm Runtime (seconds)
Source 0.23 $\pm$ 0.002
TTN 0.39 $\pm$ 0.003
TENT 3.65 $\pm$ 0.01
LAME 0.83 $\pm$ 0.03
Delta 0.92 $\pm$ 0.04
Hybrid-TTN (Uniform prior) 0.49 $\pm$ 0.01
Hybrid-TTN 0.89 $\pm$ 0.01
Table 4: Runtime of algorithms, in seconds.
A.5 Additional ablation experiments
In this section, we show ablation results for using Resnet-50 and Resnext-29 with the same configurations as used for the original Resnet-26 in CIFAR-10. Moreover, we run the three architectures using CIFAR-100 as source and CIFAR-100-C as the target domain, instead of CIFAR-10 as used in the paper. The results showed below are only for the extremely imbalanced scenario, with only one class on the target set, and for extremely corrupted cases, with severity level 5 across all corruptions.
CIFAR-10 to CIFAR-10-C
Method Resnet50 Resnet26 Resnext29
Source 60.1 $\pm$ 26.4 59.2 $\pm$ 26.5 53.1 $\pm$ 30.1
[…]
Random channel selection -20.1 $\pm$ 26.6 -14.1 $\pm$ 27.6 +3.4 $\pm$ 25.2
Hybrid-TTN (Uniform prior) -2.1 $\pm$ 24.9 -1.3 $\pm$ 22.6 +7.7 $\pm$ 19.0
Hybrid-TTN (Proposed) +1.9 $\pm$ 24.3 +3.8 $\pm$ 22.1 +8.5 $\pm$ 19.0
Hybrid-TTN (Oracle prior) +5.2 $\pm$ 24.4 +8.1 $\pm$ 22.2 +16.5 $\pm$ 20.7
CIFAR-100 to CIFAR-100-C
Method Resnet50 Resnet26 Resnext29
Source 26.1 $\pm$ 25.2 24.0 $\pm$ 25.0 25.7 $\pm$ 25.4
[…]
Random channel selection -7.9 $\pm$ 23.5 -8.4 $\pm$ 22.3 -2.8 $\pm$ 21.8
Hybrid-TTN (Uniform prior) +1.0 $\pm$ 19.3 -0.8 $\pm$ 19.9 +2.2 $\pm$ 15.8
Hybrid-TTN (Proposed) +6.1 $\pm$ 18.1 +3.4 $\pm$ 19.2 +4.5 $\pm$ 15.4
Hybrid-TTN (Oracle prior) +13.9 $\pm$ 19.9 +13.7 $\pm$ 21.0 +13.0 $\pm$ 16.5
Table 5: Ablation study using different models on CIFAR-10 and CIFAR-100. Values for Source model are accuracy, in percentages. Values for different ablated methods are the difference in accuracy between each method and the Source model. Uniform and Oracle priors are considering balanced target sets and true label distribution, respectively. Proposed method uses the pseudo-labels obtained with Uniform prior.
A.6 Detailed results
[…]